OpenStudy (anonymous):

Which of the following equalities are valid? Choose all that apply. (3²+4) - 2 =11 (3²+4) - 2 = 5 2* (3² - 1) = 16 2* (3² - 1) = 35

OpenStudy (anonymous):

The first equality is 9 +4 - 2 = 13-2 = 11. the First is true

OpenStudy (anonymous):

Which means the second is false The third is 18-2 = 16 The third is true, the fourth is false

OpenStudy (anonymous):

To recap: First and Third are True Second and Fourth are False

OpenStudy (anonymous):

Thank you so much KirbyLegs :)

OpenStudy (anonymous):

No problem. Just saving the world . . . one question at a time.